a 50 kg sled is pulled along a level surface at constant velocity by a constant force of 200n at an angle of 30° what is the coefficient of friction between the sled and the surface


1
Expert's answer
2021-12-06T09:45:07-0500

Given:

"F=200\:\rm N"

"m=50\:\rm kg"

"\theta=30^\circ"




The Newton's second law gives


"F_x=F\cos\theta-F_f=ma_x=0"

"F_y=N+F\sin\theta-mg=ma_y=0"

Also we must use the definition of friction force

"F_f=\mu N"

These three equations give

"\mu=\frac{F\cos \theta}{mg-F\sin\theta}"

"\mu=\frac{20\cos 30^\circ}{50*9.8-200\sin 30^\circ}=0.044"